Diego Company manufactures one product that is sold for $78 per unit in two geographic regions—the East and West regions. The following information pertains to the company’s first year of operations in which it produced 49,000 units and sold 44,000 units.

Variable costs per unit:
Manufacturing:
Direct materials $ 28
Direct labor $ 14
Variable manufacturing overhead $ 4
Variable selling and administrative $ 6
Fixed costs per year:
Fixed manufacturing overhead $ 686,000
Fixed selling and administrative expense $ 510,000

The company sold 32,000 units in the East region and 12,000 units in the West region. It determined that $230,000 of its fixed selling and administrative expense is traceable to the West region, $180,000 is traceable to the East region, and the remaining $100,000 is a common fixed expense. The company will continue to incur the total amount of its fixed manufacturing overhead costs as long as it continues to produce any amount of its only product.

10. What would have been the company’s variable costing net operating income (loss) if it had produced and sold 44,000 units? You do not need to perform any calculations to answer this question.
Answer #1
Variable costs per unit 52 =28+14+4+6
10
Sales units 44000
X Unit Contribution margin 26 =78-52
Contribution margin 1144000
Less: Fixed expenses 1196000 =686000+510000
Variable costing net operating loss (52000)
Variable costing net operating income will be same for both 49000 and 44000 units produced